Hello I was wondering if someone could help me with this problem. Use substitution method. Thank you
9x-2y=3
9x-6=y
The work I have done is this 9x-2y=3
3x-6=y
y=6+3x
9x-2(6+3x)=3
9x+12+6x=3
15x=3-12
15x=--9 <----- I know this isn't right, but I'm not sure how to make it right??

Answer by mananth(16946)

9x-2y=3
9x-6=y
..
plug the value of in the first equation
..
9x-2(9x-6)=3
9x-18x+12=3
-9x= -12+3
-9x=-9
x=1
plug the value of x in first equation
9-2y=3
-2y=-9+3
-2y=-6
y=3
..
CHECK
9x-2y=3
9-6 =3